multiply
(w+1/3)(w+1/4)
(w+ one over 3) (w+ one over four)

Answer by edjones(8007)

FOIL:(w+1/3)(w+1/4)
F=w^2
O=w/4
I=w/3
L=1/12
.
w^2 + 3w/12 + 4w/12 + 1/12
= w^2 + 7w/12 + 1/12
.
Ed